### v3

@@ 1.1 Notation and main results
Backward stochastic partial differential equations (BSPDEs, for short) arise in many applications of probability theory and stochastic processes. In the literature, BSPDEs can be found in various contexts, including control under incomplete information (see [Ben83, Tan98]), as adjoint equations in optimal control (see e.g. [NN90, HP91, Zho93]), in controllability problems (see e.g. [BRT03, TZ09, HSLBP23]), in mathematical finance for the formulation of the stochastic Feynman-Kac formula and the Black-Scholes equation (see [MY97, MY99]), among other applications.
@@
where $\mathcal{O}\subset\mathbb{R}^{n}$ is an open and bounded set with a $C^{2}$ boundary $\partial\mathcal{O}$. In (1), the coefficients $\alpha$ and $\beta$, the forcing term $F$ and the terminal datum $y_{T}$ are suitable random fields verifying appropriate measurability and regularity conditions.
@@
A classical question for BSPDEs consists in understanding the notion of the solution (for instance strong, mild, or weak in the PDE sense) and determining the best regularity available for the given data of the problem. This issue has been studied in the literature in different contexts and frameworks. In [HP91, LvN19], using semigroup theory, the authors establish the existence and uniqueness of mild solutions for abstract backward evolution equations in the $L^{2}$ and $L^{p}$ settings, respectively. In [HMY02, Dok12, DT12], the problem of existence and regularity of strong solutions (i.e., solutions that can be evaluated point-wisely in the spatial variable) is studied in bounded and/or unbounded domains. In the context of weak solutions (i.e., the PDE satisfies a scalar product equation), the works [DQT12, DTZ13] tackle the existence and regularity of solutions posed in the whole space within the $L^{p}$-setting and in higher order Sobolev spaces $W^{m,p}$ ($m\geq 1$, $p\geq 2$), respectively. Finally, existence and regularity of solutions in Hölder spaces are established in [TW16].
@@ 1.2 Applications
Note that we do not impose any growth or sign conditions on the nonlinear term. Thanks to the higher regularity already established for the solution, we can apply a suitable Taylor expansion to the nonlinear term, which allows us to bypass the usual difficulties associated with low regularity (see Section˜4 and Lemma˜A.3). In this way, we obtain a local well-posedness result for (7) for sufficiently small terminal data. This avoids the standard global Lipschitz assumptions commonly used in the literature (see, e.g., [HP91, HMY02] or [LZ21, Chapter 4]) and provides a more flexible framework for handling semilinear systems.
@@
Due to Theorems 1.2, 1.3 and 1.4, system (9) is well-posed for any terminal datum $y_{T}\in L^{p}_{\mathbb{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$ and any function $h\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}_{0}))$ with $p\in[2,+\infty]$. This motivates us to find controls $h$ such that the following controllability condition is fulfilled.
@@
In the $L^{2}$-setting, the (global) controllability of (11) in the sense of ˜1.8 has been studied in [HSLBP23] (see also [ZXL24]), under the assumption that $f$ is globally Lipschitz. However, as in the well-posedness results discussed earlier, this imposes a restrictive condition on the growth of $f$, which limits the applicability of the result.
@@
We say that equation (11) is locally controllable in $L^{\infty}$ if there exists $\delta>0$ such that for every terminal condition $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ with $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$, there exists a control $h\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}_{0}))$ such that the weak solution $(y,Y)$ of (11) satisfies $y(0)=0$.
@@
There is $\delta>0$ depending only on $T$, $\mathcal{O}$, and $\mathcal{O}_{0}$ such that for every $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ satisfying $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$, there exists a control $h\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}_{0}))$ such that the weak solution $(y,Y)$ of (11) satisfies $y(0)=0$.
The proof of this result combines several techniques. First, we establish a null-controllability result for the linear backward equation with an $L^{\infty}$-weighted source term, crucially relying on the cost estimate (10). To achieve this, we adapt ideas from [LTT13] in the deterministic framework and from [HSLBP22] in the stochastic setting. For completeness, we present this intermediate result in our context here. The final step of the proof employs a Banach fixed-point argument, where arguments used in Theorem˜1.7 play a key role in controlling the nonlinear terms.
@@
Theorem 1.11 provides, to the best of our knowledge, the first true local controllability result for backward semilinear SPDEs. Previous works have not directly addressed this question. The closest related result appears in Section 4 of [HSLBP22], where the authors establish a weaker notion of controllability, named statisical local controllability: instead of reaching $y(0)=0$ a.s., they prove that $\mathbb{P}(y(0)=0)\geq 1-\varepsilon$ for some $\varepsilon>0$ sufficiently small starting from a small initial data. This distinction is significant, as our result guarantees local controllability in the almost sure sense.
It is also worth noting that [HSLBP22] considers the full backward equation, including the $\beta Y$ term (compare (9) and (11)). In contrast, our approach does not treat this term because of the limitations in the linear controllability analysis (see Proposition˜1.9). Nevertheless, it does establish a controllability result in a setting where previously only high-probability statements were known.
@@
Theorem˜1.11 does not cover the case in which the source term $f$ depends on $Y$. Indeed, in the key a priori estimate of the source-term method in Proposition˜5.1, the source term is required to lie in $L^{\infty}$, while our analysis only yields $L^{p}$-regularity for the process $Y$ (with finite $p$), which is generally insufficient to close the estimates. This obstruction already appears at the level of $L^{2}$, as noted in our previous work (see [HSLBP22, Remark 4.4]), where the source-term method does not provide suitable bounds when $f$ depends on $Y$. Addressing such nonlinear dependencies would likely require methods of a different nature, and we view this as an interesting direction for future research.
@@
$\begin{cases}\textnormal{d}y=(\Delta y+\chi_{\mathcal{O}_{0}}h+f(y))\,\textnormal{d}t+(ay+g(y))\,\textnormal{d}W(t)&\text{in }(0,T)\times\mathcal{O},\\ y=0&\text{on }(0,T)\times\Gamma,\\ y(0)=y_{0}&\text{in }\mathcal{O},\end{cases}$ (12)
@@
Let us recall that the linear case (i.e. $f\equiv 0$, $g\equiv 0$) has been studied extensively. Several controllability results are available under different assumptions, relying on Carleman estimates (see [FI96]) and on the Lebeau–Robbiano strategy [LR95], with both approaches requiring suitable adaptations to the stochastic setting (see, e.g., [BRT03, TZ09, L1̈1, Liu14]). We also point out that, in some of these works, an additional control acting in the diffusion term is required; namely, the term $ay\,\textnormal{d}W(t)$ in (12) is replaced by $(ay+H)\,\textnormal{d}W(t)$ for a suitably chosen control $H$.
When nonlinearities are present, particularly in the diffusion term, the analysis becomes substantially more delicate. In the globally Lipschitz setting, our previous work [HSLBP23] (see also [ZXL24]) established global controllability but—as in the linear case—only by introducing an additional control in the diffusion term. Subsequently, in [HSLBP22], we proved statistical local controllability with a single control (that is, $\mathbb{P}(y(T)=0)\geq 1-\varepsilon$ for $\varepsilon>0$ small and small initial data), under suitable structural assumptions on the nonlinear diffusion term $g(y)$, essentially in one spatial dimension. However, the general problem of proving true local controllability for semilinear equations with nonlinear noise remains open.
@@ 2 Itô’s formula for the $L^{p}$-norm
The goal of this section is to state the Itô’s formula for the $L^{p}$-norm and deduce $L^{p}$-estimates for the weak solution $(y,Y)$ of (1). Note that during this section, the $L^{p}$-estimates are not uniform with respect to $p$. The final conclusion of this section will be the proof of Theorem˜1.3.
The goal of this section is to prove the following result, which is essential for the proofs of Theorem˜1.3 and Theorem˜1.4. This result includes an initial a priori estimate and an Itô’s formula for the $L^{p}$-norm of the solution to the backward SPDE (1). More precisely, we have the following.
@@
$=\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}y\left(\alpha y+\beta Y+F\right)\,\textnormal{d}x-p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}yYdW(s),\quad\textnormal{a.s.}$ (14)
@@ 2.2 Proof of the Ito’s formula for the $L^{p}$-norm
for all $t\in[0,T]$. We will see that each term in (38) converges to the corresponding one in (14). At this point, we will make use of estimate (13).
@@
$\quad+C\left(\int_{0}^{T}\int_{\mathcal{O}}|yF|\,\textnormal{d}x\textnormal{d}{s}+\left|\int_{t}^{T}\int_{\mathcal{O}}yY\,\textnormal{d}x\textnormal{d}{W}(s)\right|^{p/2}\right),$ (42)
@@
$\left(\int_{0}^{T}\int_{\mathcal{O}}|yF|\,\textnormal{d}x\textnormal{d}t\right)^{p/2}\leq\sup_{s\in[0,T]}\|y(s)\|^{p/2}_{L^{2}(\mathcal{O})}\left(\int_{0}^{T}\|F(s)\|_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p/2},\quad\textnormal{a.s.}$
@@
${\mathbb{E}}\left[\left(\int_{0}^{T}\int_{\mathcal{O}}|yF|\,\textnormal{d}x\textnormal{d}t\right)^{p/2}\right]\leq\frac{1}{2}{\mathbb{E}}\left(\sup_{s\in[0,T]}\|y(s)\|^{p}_{L^{2}(\mathcal{O})}\right)+\frac{1}{2}{\mathbb{E}}\left[\left(\int_{0}^{T}\|F(s)\|_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p}\right].$ (45)
@@ 3 Uniform $L^{p}$- and $L^{\infty}$-estimates
Now, we take $r=t$ in the previous estimate and use the fact that $\mathbb{E}(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\mid\mathcal{F}_{t})=\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}$ since $y$ if $\mathcal{F}_{t}$-adapted to obtain
@@
${\mathbb{E}}\left(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\right)\leq\exp([pK+(p-1)](T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}(\Omega;L^{p}(0,T;L^{p}(\mathcal{O}))}^{p}\right),$
@@
$\left\lVert y(t)\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}^{p}\leq\exp([pK+(p-1)](T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}(\Omega;L^{p}(0,T;L^{p}(\mathcal{O}))}^{p}\right).$
@@
$\left\lVert y(t)\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}$ $\leq\exp((K+(p-1)/p)(T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}(\Omega;L^{p}(0,T;L^{p}(\mathcal{O}))}^{p}\right)^{1/p}$
$\leq\exp([K+1](T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}+\left\lVert F\right\rVert_{L^{p}(\Omega;L^{p}(0,T;L^{p}(\mathcal{O}))}\right),$

### v5

@@ 1.1 Notation and main results
Backward stochastic partial differential equations (BSPDEs, for short) arise in many applications of probability theory and stochastic processes. In the literature, BSPDEs can be found in various contexts, including control under incomplete information (see [BEN83, TAN98]), as adjoint equations in optimal control (see e.g. [NN90, HP91, ZHO93]), in controllability problems (see e.g. [BRT03, TZ09, HLP23]), in mathematical finance for the formulation of the stochastic Feynman-Kac formula and the Black-Scholes equation (see [MY97, MY99]), among other applications.
@@
where $\mathcal{O}\subset\mathbb{R}^{d}$ ($d\geq 1$) is an open and bounded set with a $C^{2}$ boundary $\partial\mathcal{O}$. In (1), the coefficients $\alpha$ and $\beta$, the forcing term $F$ and the terminal datum $y_{T}$ are suitable random fields verifying appropriate measurability and regularity conditions.
@@
A classical question for BSPDEs consists in understanding the notion of the solution (for instance strong, mild, or weak in the PDE sense) and determining the best regularity available for the given data of the problem. This issue has been studied in the literature in different contexts and frameworks. In [HP91, Lv01], using semigroup theory, the authors establish the existence and uniqueness of mild solutions for abstract backward evolution equations in the $L^{2}$ and $L^{p}$ settings, respectively. In [HMY02, DOK12, DT12], the problem of existence and regularity of strong solutions (i.e., solutions that can be evaluated point-wisely in the spatial variable) is studied in bounded and/or unbounded domains. In the context of weak solutions (i.e., the PDE satisfies a scalar product equation), the works [DQT12, DTZ13] tackle the existence and regularity of solutions posed in the whole space within the $L^{p}$-setting and in higher order Sobolev spaces $W^{m,p}$ ($m\geq 1$, $p\geq 2$), respectively. Finally, existence and regularity of solutions in Hölder spaces are established in [TW16].
@@ 1.2 Applications
Note that we do not impose any growth or sign conditions on the nonlinear term. Thanks to the higher regularity already established for the solution, we can apply a suitable Taylor expansion to the nonlinear term, which allows us to bypass the usual difficulties associated with low regularity (see Section˜4 and Lemma˜A.3). In this way, we obtain a local well-posedness result for (7) for sufficiently small terminal data. This avoids the standard global Lipschitz assumptions commonly used in the literature (see, e.g., [HP91, HMY02] or [LZ02, Chapter 4]) and provides a more flexible framework for handling semilinear systems.
@@
Due to Theorems 1.2, 1.3 and 1.4, system (9) is well-posed for any terminal datum $y_{T}\in L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))$ and any function $h\in L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}_{0}))$ with $p\in[2,+\infty]$. This motivates us to find controls $h$ such that the following controllability condition is fulfilled.
@@
In the $L^{2}$-setting, the (global) controllability of (11) in the sense of ˜1.8 has been studied in [HLP23] (see also [ZXL24]), under the assumption that $f$ is globally Lipschitz. However, as in the well-posedness results discussed earlier, this imposes a restrictive condition on the growth of $f$, which limits the applicability of the result.
@@
We say that equation (11) is locally controllable in $L^{\infty}$ if there exists $\delta>0$ such that for every terminal condition $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ with $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$, there exists a control $h\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}_{0}))$ such that the weak solution $(y,Y)$ of (11) satisfies $y(0)=0$ a.s.
@@
There is $\delta>0$ depending only on $T$, $\mathcal{O}$, and $\mathcal{O}_{0}$ such that for every $y_{T}\in L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))$ satisfying $\|y_{T}\|_{L^{\infty}_{\mathcal{F}_{T}}(\Omega;L^{\infty}(\mathcal{O}))}\leq\delta$, there exists a control $h\in L^{\infty}_{\mathbb{F}}(0,T;L^{\infty}(\mathcal{O}_{0}))$ such that the weak solution $(y,Y)$ of (11) satisfies $y(0)=0$ a.s.
The proof of this result combines several techniques. First, we establish a null-controllability result for the linear backward equation with an $L^{\infty}$-weighted source term, crucially relying on the cost estimate (10). To achieve this, we adapt ideas from [LTT13] in the deterministic framework and from [HLP22] in the stochastic setting. For completeness, we present this intermediate result in our context here. The final step of the proof employs a Banach fixed-point argument, where arguments used in Theorem˜1.7 play a key role in controlling the nonlinear terms.
@@
Theorem 1.11 provides, to the best of our knowledge, the first true local controllability result for backward semilinear SPDEs. Previous works have not directly addressed this question. The closest related result appears in Section 4 of [HLP22], where the authors establish a weaker notion of controllability, named statisical local controllability: instead of reaching $y(0)=0$ a.s., they prove that $\mathbb{P}(y(0)=0)\geq 1-\varepsilon$ for some $\varepsilon>0$ sufficiently small starting from a small initial data. This distinction is significant, as our result guarantees local controllability in the almost sure sense.
It is also worth noting that [HLP22] considers the full backward equation, including the $\beta Y$ term (compare (9) and (11)). In contrast, our approach does not treat this term because of the limitations in the linear controllability analysis (see Proposition˜1.9). Nevertheless, it does establish a controllability result in a setting where previously only high-probability statements were known.
@@
Theorem˜1.11 does not cover the case in which the source term $f$ depends on $Y$. Indeed, in the key a priori estimate of the source-term method in Proposition˜5.1, the source term is required to lie in $L^{\infty}$, while our analysis only yields $L^{p}$-regularity for the process $Y$ (with finite $p$), which is generally insufficient to close the estimates. This obstruction already appears at the level of $L^{2}$, as noted in our previous work (see [HLP22, Remark 4.4]), where the source-term method does not provide suitable bounds when $f$ depends on $Y$. Addressing such nonlinear dependencies would likely require methods of a different nature, and we view this as an interesting direction for future research.
@@
$\begin{cases}\textnormal{d}y=(\Delta y+\chi_{\mathcal{O}_{0}}h+f(y))\,\textnormal{d}t+(ay+g(y))\,\textnormal{d}W(t)&\text{in }(0,T)\times\mathcal{O},\\ y=0&\text{on }(0,T)\times\partial\mathcal{O},\\ y(0)=y_{0}&\text{in }\mathcal{O},\end{cases}$ (12)
@@
Let us recall that the linear case (i.e. $f\equiv 0$, $g\equiv 0$) has been studied extensively. Several controllability results are available under different assumptions, relying on Carleman estimates (see [FI96]) and on the Lebeau–Robbiano strategy [LR95], with both approaches requiring suitable adaptations to the stochastic setting (see, e.g., [BRT03, TZ09, LÜ11, LIU14]). We also point out that, in some of these works, an additional control acting in the diffusion term is required; namely, the term $ay\,\textnormal{d}W(t)$ in (12) is replaced by $(ay+H)\,\textnormal{d}W(t)$ for a suitably chosen control $H$.
When nonlinearities are present, particularly in the diffusion term, the analysis becomes substantially more delicate. In the globally Lipschitz setting, our previous work [HLP23] (see also [ZXL24]) established global controllability but—as in the linear case—only by introducing an additional control in the diffusion term. Subsequently, in [HLP22], we proved statistical local controllability with a single control (that is, $\mathbb{P}(y(T)=0)\geq 1-\varepsilon$ for $\varepsilon>0$ small and small initial data), under suitable structural assumptions on the nonlinear diffusion term $g(y)$, essentially in one spatial dimension. However, the general problem of proving true local controllability for semilinear equations with nonlinear noise remains open.
@@ 2 Itô’s formula for the $L^{p}$-norm
The goal of this section is to state Itô’s formula for the $L^{p}$-norm and deduce $L^{p}$-estimates for the weak solution $(y,Y)$ of (1). Note that throughout this section, the $L^{p}$-estimates are not uniform with respect to $p$. The final conclusion will be the proof of Theorem˜1.3.
We begin by proving the following result, which is essential for the proofs of Theorem˜1.3 and Theorem˜1.4. It provides an initial a priori estimate together with an Itô’s formula for the $L^{p}$-norm of the solution to the backward SPDE (1). More precisely, we have the following.
@@
$=\int_{\mathcal{O}}|y_{T}|^{p}\,\textnormal{d}x+p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}y\left(\alpha y+\beta Y+F\right)\,\textnormal{d}x\textnormal{d}{s}-p\int_{t}^{T}\int_{\mathcal{O}}|y|^{p-2}yY\textnormal{d}{x}dW(s),\quad\textnormal{a.s.}$ (14)
@@ 2.2 Proof of the Ito’s formula for the $L^{p}$-norm
for all $t\in[0,T]$. We will see that each term in (38) converges to the corresponding one in (14) as $n\to\infty$. At this point, we will make use of estimate (13).
@@
$\quad+C\left(\left(\int_{0}^{T}\int_{\mathcal{O}}|yF|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}+\left|\int_{t}^{T}\int_{\mathcal{O}}yY\,\textnormal{d}x\textnormal{d}{W}(s)\right|^{p/2}\right),$ (42)
@@
$\left(\int_{0}^{T}\int_{\mathcal{O}}|yF|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}\leq\sup_{s\in[0,T]}\|y(s)\|^{p/2}_{L^{2}(\mathcal{O})}\left(\int_{0}^{T}\|F(s)\|_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p/2},\quad\textnormal{a.s.}$
@@
${\mathbb{E}}\left[\left(\int_{0}^{T}\int_{\mathcal{O}}|yF|\,\textnormal{d}x\textnormal{d}{s}\right)^{p/2}\right]\leq\frac{1}{2}{\mathbb{E}}\left(\sup_{s\in[0,T]}\|y(s)\|^{p}_{L^{2}(\mathcal{O})}\right)+\frac{1}{2}{\mathbb{E}}\left[\left(\int_{0}^{T}\|F(s)\|_{L^{2}(\mathcal{O})}\textnormal{d}{s}\right)^{p}\right].$ (45)
@@ 3 Uniform $L^{p}$- and $L^{\infty}$-estimates
Now, we take $r=t$ in the previous estimate and use the fact that $\mathbb{E}(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\mid\mathcal{F}_{t})=\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}$ since $y$ is $\mathcal{F}_{t}$-adapted to obtain
@@
${\mathbb{E}}\left(\left\lVert y(t)\right\rVert_{L^{p}(\mathcal{O})}^{p}\right)\leq\exp([pK+(p-1)](T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}^{p}\right),$
@@
$\left\lVert y(t)\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}^{p}\leq\exp([pK+(p-1)](T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}^{p}\right).$
@@
$\left\lVert y(t)\right\rVert_{L^{p}(\Omega;L^{p}(\mathcal{O}))}$ $\leq\exp((K+(p-1)/p)(T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}^{p}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}^{p}\right)^{1/p}$
$\leq\exp([K+1](T-t))\left(\left\lVert y_{T}\right\rVert_{L^{p}_{\mathcal{F}_{T}}(\Omega;L^{p}(\mathcal{O}))}+\left\lVert F\right\rVert_{L^{p}_{\mathbb{F}}(0,T;L^{p}(\mathcal{O}))}\right),$